### v3

@@ 4.1. Maps between sets of tropical points
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair. We extend $\widetilde{B}_{t_{0}}$ to an $m\times m$ skew-symmetrizable matrix $\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}}=(\widetilde{B}_{t_{0}}\mid M)$, where $M$ is an $m\times(m-n)$ matrix. Denote by $\{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}\mid t\in\mathbb{T}_{n}\}$ the collection of matrices obtained from $\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}}$ by any sequence of mutations in directions $1,\ldots,n$. Let $\widetilde{S}=diag(s_{1},\ldots,s_{m})$ be a skew-symmetrizer of the transposed matrix $(\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}})^{T}$. In particular, $\widetilde{S}(\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}})^{T}$ is a skew-symmetric matrix.
@@
Keep the above setting. Then
$[{\bf a}]:=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\mapsto\{\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
defines a map from the tropical points in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$.
Proof.
Denote by ${\bf g}^{t}:=\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. We will use Lemma 4.1 to show the collection $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$.
Let ${t}$${t^{\prime}}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. Since $[{\bf a}]=\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ and by Lemma 4.2, we have ${\bf a}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$, where
$\varepsilon_{0}=\begin{cases}-,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0;\\ +,&\text{if }\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}<0.\end{cases}$
By (2.2), we have $\widetilde{B}_{t^{\prime}}^{\operatorname{\mathsf{sq}}}=E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}}\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}}$. Thus $(\widetilde{B}_{t^{\prime}}^{\operatorname{\mathsf{sq}}})^{T}=(F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}$. Since $\widetilde{S}$ is a skew-symmetrizer of $(\widetilde{B}_{t_{0}}^{\operatorname{\mathsf{sq}}})^{T}$, we have
$\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}=-\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}\widetilde{S}$
for any vertex $t\in\mathbb{T}_{n}$. In the following proofs, we will use the following obvious relations:
$\widetilde{S}(F_{k,\pm}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}=(E_{k,\pm}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})\widetilde{S},\;\;\;E_{k,\pm}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}}=E_{k,\pm}^{\widetilde{B}_{t}},\;\;\;\text{and}\;\;\;(E_{k,\pm}^{\widetilde{B}_{t}})^{2}=I_{m}.$
Now we have
${\bf g}^{t^{\prime}}$ $=$ $\widetilde{S}(\widetilde{B}_{t^{\prime}}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t^{\prime}}=\widetilde{S}\left((F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}\right)\cdot(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}{\bf a}^{t}$
$=$ $\widetilde{S}(F_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})^{T}\cdot(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}\cdot(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}})^{T}\cdot{\bf a}^{t}$
$=$ $(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}}})\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}$
$=$ $(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}}){\bf g}^{t}.$
Now we prove $\varepsilon_{0}=\varepsilon(g_{k}^{t})$, where $\varepsilon(g_{k}^{t})$ is as given in Lemma 4.1. By the definitions of $\varepsilon(g_{k}^{t})$ and $\varepsilon_{0}$, it suffices to show that $g_{k}^{t}\geq 0$ if and only if $\sum_{j=1}^{m}b_{jk;t}a_{j}^{t}\geq 0$. This is clear, because ${\bf g}^{t}=\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}$ implies that $g_{k}^{t}=s_{k}\cdot(\sum_{j=1}^{m}b_{jk;t}a_{j}^{t})$.
Hence, ${\bf g}^{t^{\prime}}=(E_{k,\varepsilon_{0}}^{\widetilde{B}_{t}}){\bf g}^{t}=(E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}){\bf g}^{t}$. By Lemma 4.1, we know the collection
$\{{\bf g}^{t}=\widetilde{S}(\widetilde{B}_{t}^{\operatorname{\mathsf{sq}}})^{T}{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. This completes the proof. ∎
Proposition 4.4.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$ be a Langlands-Poisson triple. Then
$[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\mapsto\{\Lambda_{t}{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
@@
Denote by ${\bf a}^{t}:=\Lambda_{t}{\bf g}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. We will use Lemma 4.2 to show the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$.
Let ${t}$${t^{\prime}}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. Since $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ and by Lemma 4.1, we have ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$, where
@@ 4.2. Tropical invariant and $F$-invariant
$F[{\bf r}]:=\max\{{\bf v}^{T}{\bf r}\mid c_{{\bf v}}\neq 0\}\in\mathbb{Z}.$
We call the map $F[-]:\mathbb{Z}^{n}\rightarrow\mathbb{Z}$ a tropical polynomial.
@@
    If $0\neq F\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ has constant term $1$, then $F[{\bf r}]\geq 0$ for any ${\bf r}\in\mathbb{Z}^{n}$.
@@
Now we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$, and let $\mathcal{U}$ be the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where ${\bf 0}$ is the $n\times(m-n)$ zero matrix. By Proposition 2.2, the mutation matrices of $\mathcal{U}$ always have full rank.
We denote by $\mathcal{U}^{[{\bf g}]}$ the set of $[{\bf g}]$-good elements in $\mathcal{U}$ and by
$\mathcal{U}^{\operatorname{\mathsf{good}}}:=\bigcup_{[{\bf g}]\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})}\mathcal{U}^{[\bf g]}$
Definition 4.7 (Tropical pairing).
Let $u,u^{\prime}\in\mathcal{U}^{\operatorname{\mathsf{good}}}$ and consider their canonical expressions with respect to a vertex $w\in\mathbb{T}_{n}$:
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),\;\;\;u^{\prime}={\bf x}_{w}^{{\bf g}_{u^{\prime}}^{w}}F_{u^{\prime}}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}).$
Define an integer $\langle u,u^{\prime}\rangle_{w}$ as follows:
$\langle u,u^{\prime}\rangle_{w}:=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}].$
The pairing $\langle-,-\rangle_{w}:\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}$ is called the tropical pairing at vertex $w\in\mathbb{T}_{n}$.
@@
For any two good elements $u$ and $u^{\prime}$, the integer $\langle u,u^{\prime}\rangle_{w}$ only depends on $u$ and $u^{\prime}$, not on the choice of $w\in\mathbb{T}_{n}$. We call $\langle u,u^{\prime}\rangle_{w}$ the tropical invariant between $u$ and $u^{\prime}$.
@@
For a vertex $t\in\mathbb{T}_{n}$, let $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})$ and $u^{\prime}={\bf x}_{t}^{{\bf g}_{u^{\prime}}^{t}}F_{u^{\prime}}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t})$ be the canonical expressions of $u$ and $u^{\prime}$ with respect to $t\in\mathbb{T}_{n}$. We know that
$[{\bf g}_{u^{\prime}}]:=\{{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. By Proposition 4.4, the collection
$[{\bf a}]:=\{{\bf a}^{t}=\Lambda_{t}{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. By Corollary 3.5, the tropical point $[{\bf a}]\in\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism
$\beta_{u^{\prime}}:\mathbb{F}_{>0}:=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that $\beta_{u^{\prime}}({\bf x}_{t})=({\bf a}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. Notice that the simifield homomorphism $\beta_{u^{\prime}}$ is uniquely determined by $u^{\prime}$.

### v4

@@ 4.1. Semifield homomorphisms arising from good elements
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ be a Langlands-Poisson triple. Then
${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}\mapsto\{\Lambda_{t}{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
@@
Denote by ${\bf a}^{t}\coloneqq\Lambda_{t}{\bf g}^{t}=(a_{1}^{t},\ldots,a_{m}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. We will use Lemma 4.2 to show the collection $\{{\bf a}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$.
Let ${t}$${t^{\prime}}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. Since ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ and by Lemma 4.1, we have ${\bf g}^{t^{\prime}}=E_{k,\varepsilon(g_{k}^{t})}^{\widetilde{B}_{t}}{\bf g}^{t}$, where
@@
Proposition 4.4.
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ be a Langlands-Poisson triple and let $\mathcal{U}$ be the corresponding $\mathsf{\Lambda}$-upper cluster algebra. Then any good element $u\in\mathcal{U}$ defines a unique semifield homomorphism
$\beta_{u}:\;\mathbb{F}_{>0}:=\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that $\beta_{u}({\bf x}_{t})=(\Lambda_{t}{\bf g}_{u}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$, where ${\bf g}_{u}^{t}\in\mathbb{Z}^{m}$ is the extended $g$-vector of $u$ with respect to vertex $t$.
Proof.
Since $u$ is a good element, we know that the collection
${\bf g}_{u}]\coloneqq\{{\bf g}_{u}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. Then by Proposition 4.3, the collection
${\bf a}]\coloneqq\{{\bf a}^{t}=\Lambda_{t}{\bf g}_{u}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$
is a tropical point in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. By Corollary 3.5, the tropical point ${\bf a}]\in\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ corresponds to a unique semifield homomorphism
$\beta_{u}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$
such that $\beta_{u}({\bf x}_{t})=({\bf a}^{t})^{T}=(\Lambda_{t}{\bf g}_{u}^{t})^{T}$ for any vertex $t\in\mathbb{T}_{n}$. ∎
Remark 4.5.
We remark that the above proof still works if we replace the good elements by compatibly pointed elements. But in this paper, we only use good elements.
@@ 4.2. Tropical invariant and $F$-invariant
In this subsection, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$, and let $\mathcal{U}$ be the corresponding $\mathsf{\Lambda}$-upper cluster algebra. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where $\bf 0$} is the $n\times(m-n)$ zero matrix. By Proposition 2.2, the mutation matrices of $\mathcal{U}$ always have full rank.
We denote by $\mathcal{U}^{[{\bf g}]}$ the set of $\bf g$}]-good elements in $\mathcal{U}$ and by
$\mathcal{U}^{\operatorname{\mathsf{good}}}\coloneqq\bigcup_{[{\bf g}]\in\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})}\mathcal{U}^{[\bf g]}$
Definition 4.6 (Tropical invariant and $F$-invariant).
Let $\mathcal{U}$ be a $\mathsf{\Lambda}$-upper cluster algebra.
(i) For $u,u^{\prime}\in\mathcal{U}^{\operatorname{\mathsf{good}}}$, the tropical invariant $\langle u,u^{\prime}\rangle_{\rm trop}$ of the pair $(u,u^{\prime})$ is defined by
$\langle u,u^{\prime}\rangle_{\rm trop}\coloneqq\beta_{u^{\prime}}(u)\;\in\mathbb{Z},$
where $\beta_{u^{\prime}}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$ is the semifield homomorphism defined in Proposition 4.4. The pairing $\langle-,-\rangle_{\rm trop}\colon\mathcal{U}^{\operatorname{\mathsf{good}}}\times\mathcal{U}^{\operatorname{\mathsf{good}}}\rightarrow\mathbb{Z}$ is called the tropical pairing on the $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$.
(ii) For $u,u^{\prime}\in\mathcal{U}^{\operatorname{\mathsf{good}}}$, the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ of the pair $(u,u^{\prime})$ is defined by
$(u\mid\mid u^{\prime})_{F}\coloneqq\beta_{u^{\prime}}(u)+\beta_{u}(u^{\prime})=\langle u,u^{\prime}\rangle_{\rm trop}+\langle u^{\prime},u\rangle_{\rm trop}.$
In the case $(u\mid\mid u^{\prime})_{F}=0$, we say that $u$ and $u^{\prime}$ are $F$-compatible.
Remark 4.7.
Notice that in general $\langle u,u^{\prime}\rangle_{\rm trop}\neq\langle u^{\prime},u\rangle_{\rm trop}$, but we always have $(u\mid\mid u^{\prime})_{F}=(u^{\prime}\mid\mid u)_{F}$, by definition.
@@
$F[{\bf r}]\coloneqq\max\{{\bf v}^{T}{\bf r}\mid c_{{\bf v}}\neq 0\}\in\mathbb{Z}.$
We call the map $F[-]\colon\mathbb{Z}^{n}\rightarrow\mathbb{Z}$ a tropical polynomial.
@@
    If $F\in\mathbb{Z}[y_{1},\ldots,y_{n}]$ has constant term $1$, then $F[{\bf r}]\geq 0$ for any ${\bf r}\in\mathbb{Z}^{n}$.
@@
Proof.
(i) This is clear. (ii) Fix ${\bf r}=(r_{1},\ldots,r_{n})^{T}\in\mathbb{Z}^{n}$ and let
$\pi_{\bf r}\colon\mathbb{Q}_{\rm sf}(y_{1},\ldots,y_{n})\rightarrow\mathbb{Z}^{\max}$
be the semifield homomorphism induced by $y_{i}\mapsto r_{i}$. For any $0\neq F\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, we have $F\in\mathbb{Q}_{\rm sf}(y_{1},\ldots,y_{n})$ and $\pi_{\bf r}(F)=F[{\bf r}]$. Since $0\neq F_{1},F_{2}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$, we have $\pi_{\bf r}(F_{1}F_{2})=\pi_{\bf r}(F_{1})+\pi_{\bf r}(F_{2})$, i.e., $(F_{1}F_{2})[{\bf r}]=F_{1}[{\bf r}]+F_{2}[{\bf r}]$. ∎
The following result shows that for each $\mathsf{\Lambda}$-seed $({\bf x}_{w},\widetilde{B}_{w},\Lambda_{w})$ of $\mathcal{U}$, we have an explicit formula to calculate the tropical invariant and $F$-invariant.
@@
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. Let
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),\;\;\;u^{\prime}={\bf x}_{w}^{{\bf g}_{u^{\prime}}^{w}}F_{u^{\prime}}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$
be the canonical expressions of $u$ and $u^{\prime}$ with respect to any vertex $w\in\mathbb{T}_{n}$. Then we have
  • (i)
    $\langle u,u^{\prime}\rangle_{\rm trop}=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$.
  • (ii)
    $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]$. In particular, $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
  • (iii)
    $(u\mid\mid u)_{F}=0$ whenever $u$ is a cluster monomial.
@@
(i) By definition, $\langle u,u^{\prime}\rangle_{\rm trop}=\beta_{u^{\prime}}(u)$, where $\beta_{u^{\prime}}$ is the unique semifield homomorphism
$\beta_{u^{\prime}}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max}$